OpenStudy (nhra34x0):

8) Kara went to the store to buy some vegetables. She fills her grocery bag with celery and radishes. The total weight is 2 1/2 lbs. The celery cost $2.50 per pound and the radishes cost $4.80 per pound. She spent $10.85. How many pounds of each vegetable did she buy?

OpenStudy (3mar):

May I help?

OpenStudy (nhra34x0):

yes please. Im so lost

OpenStudy (mikey):

she bought two pounds of radishes and a pound of celery 4.8+4.8+2.5

OpenStudy (cuanchi):

NO, the total weight is 2 1/2Lb you have to design a system of two equations with two unknowns (x,y)

OpenStudy (mikey):

oh I didn't see that part

OpenStudy (mikey):

my bad don't listen to me

OpenStudy (cuanchi):

x= lb celery y= lb radish x + y = 1.5 lb 2.50 x + 4.80y = 10.85

OpenStudy (nhra34x0):

Where did you get 1.5 ?

OpenStudy (3mar):

He meant that x+y=2.5 ib as the total amount of type types is 2.5 ib Can you solve these two equations together? I think it it easy!

OpenStudy (nhra34x0):

I already got it haha, I got notifcations about it and was wondering where 1.5 came from. Thank you guys though!

OpenStudy (cuanchi):

yes sorry 2.5 that was my fault now

OpenStudy (3mar):

No problem. We all do. What matters is he gets the idea!